Divide write the answer in scientific notation.

85.5x10 to the negative 7th power by 9.5x10 to the negative 4th power

Answer by checkley71(8403)

85.5*10^-7/9.5*10^-4
88.5/9.5=9 for the integer component.
10^-7=1/10^7
1/10^-4=10^4
combining these elements we get:
10^4/10^7=1/10^3
thus the answer is:
9/10^3=9*10^-3 answer.